# Multipartite entanglement distribution in a topological photonic network

### v1

@@ Appendix B Details on the control pulses
The pulses used have three parts: an initial preparation time where the control parameter is switched on slowly during an interval of $t_{\mathrm{prep}}$, a time where the parameter is kept constant, and a final interval lasting $t_{\mathrm{prep}}$ where it is switched off slowly. All pulses used are symmetric, and $t_{\mathrm{prep}}$ is chosen so that $t_{\mathrm{prep}}\gg w^{-1}$, in order for the process to be adiabatic with respect to the energy gap between bulk and edge states. The same kind of pulses, although in different models, were used in Zurita et al. (2023).
@@
$v(t)=\begin{cases}v_{0}\sin^{2}(\Omega t)&\textrm{if }0\leq t\leq t_{\mathrm{prep}}\\ v_{0}&\textrm{if }t_{\mathrm{prep}}<0\leq t_{\mathrm{tr}}-t_{\mathrm{prep}}\\ v_{0}\sin^{2}[\Omega(t-t_{\mathrm{tr}})]&\textrm{if }t_{\mathrm{tr}}-t_{\mathrm{prep}}<t\leq t_{\mathrm{tr}},\end{cases}$ (14)
@@
When the number of domains is $n_{d}>2$, the control parameters need to be different in different domains for the transfer time to be minimal. The height of the pulses are chosen symmetrically with respect to the center of the transfer length, and they are found numerically once the $v_{\mathrm{tr}}$ (the height at the first and last domains) is fixed. This is explained in depth in Zurita et al. (2023).
@@
We include below (Table 1) the parameters for all protocols used in this work, when applicable: domain number $n_{d}$, domain length $\ell$, sites involved in each transfer $L$, control parameter heights $v_{\mathrm{tr}}$ and $v_{\mathrm{tr}}^{\prime}$ (for the central domains in the 4-domain case), barrier height $v_{\mathrm{bar}}$, and the times for preparation ($t_{\mathrm{prep}}$), barrier preparation ($t_{\mathrm{bar}}$) and total duration ($t_{\mathrm{tr}}$).
[…]
@@
Table 1: Parameters for all protocols used, and the states prepared with it.
@@ Appendix C Acquired phases and phase-shift protocols
$|\mathcal{P}_{k}\rangle$ $=\mathcal{N}_{\mathcal{P}}\left[\sum_{j=j_{0}-\ell/2}^{j_{0}-1}\left(-\frac{w}{v_{l}}\right)^{j-j_{0}+1}|j,\alpha(k)\rangle\right.$
$\left.-\sum_{j=j_{0}+1}^{j_{0}+\ell/2}\left(-\frac{w}{v_{r}}\right)^{j_{0}-j+1}|j,\alpha(k)\rangle\right]$ (17)
$|\mathcal{R}\rangle$ $=-\mathcal{N}_{\mathcal{R}}\sum_{j=N_{c}-\ell/2}^{N_{c}}\left(-\frac{w}{v}\right)^{-j+N_{c}+1}|j,\alpha(N)\rangle$ (18)
@@
Appendix D Fidelity and bulk-edge decoupling
@@
where $|\psi_{0}\rangle$ is the intended final state, and $\rho_{Q}$ is the density matrix of the qudit system at the end of the realization, after tracing out the photonic degrees of freedom. For every protocol considered in this work, we obtain a fidelity higher than 0.99 for the pristine system, which is the current error-correction threshold Fowler et al. (2012). We include the $\Phi$, W and GHZ states prepared as in the main text, and the $\Psi$ state, prepared as detailed in Appendix E.
@@
where $\rho_{ph}$ is the density matrix of the photonic degrees of freedom. The probability $p(\textrm{no bulk})$ is then found as $p(\textrm{no bulk})=p(\varnothing_{\gamma})+p(\textrm{edge})$. We can see that it is extremely close to 1, even in the presence of general disorder. This shows that the bulk-edge decoupling caused by the topology of the system is remarkably robust. This has been reported for other topological systems in the literature Lang and Büchler (2017). The three-qudit cases show a slightly smaller decoupling due to the use of trivial domains as barriers, which can allow some photon amplitude to leak into the bulk. This problem could be mitigated by decoupling the domains altogether by switching off some $w$ amplitudes, but we decided to keep them constant throughout in this work for simplicity.
@@
Appendix E GHZ-like protocol for Bell states
@@
Appendix F Generalization to any number of qubits
@@
where $\zeta_{k,k^{\prime}}$ is the phase acquired in the transfer from the cavity housing qubit $k$ to the cavity housing qubit $k^{\prime}$, which can be obtained using Eq. 19.
@@
The value of $\eta$ in the initial state can be chosen to fix the relative phase. Both protocols can be easily modified via the qudit-cavity coupling operations to change the relative weights between their final components, and the relative phases in the W state can be adjusted with phase gates or using the phase-shift protocols in Appendix C.

### v2

@@ Appendix B Details on the control pulses
The pulses used have three parts: an initial preparation time where the control parameter is switched on slowly during an interval of $t_{\mathrm{prep}}$, a time where the parameter is kept constant, and a final interval lasting $t_{\mathrm{prep}}$ where it is switched off slowly. All pulses used are symmetric, and $t_{\mathrm{prep}}$ is chosen so that $t_{\mathrm{prep}}\gg w^{-1}$, in order for the process to be adiabatic with respect to the energy gap between bulk and edge states. The same kind of pulses, although in different models, were used in [19]. We keep $w=1$ throughout.
@@
$v(t)=\begin{cases}v_{0}\sin^{2}(\Omega t)&\textrm{if }0\leq t\leq t_{\mathrm{prep}}\\ v_{0}&\textrm{if }t_{\mathrm{prep}}<0\leq t_{\mathrm{tr}}-t_{\mathrm{prep}}\\ v_{0}\sin^{2}[\Omega(t-t_{\mathrm{tr}})]&\textrm{if }t_{\mathrm{tr}}-t_{\mathrm{prep}}<t\leq t_{\mathrm{tr}},\end{cases}$
@@
When the number of domains is $n_{d}>2$, the control parameters need to be different in different domains for the transfer time to be minimal. The height of the pulses are chosen symmetrically with respect to the center of the transfer length, and they are found numerically once the $v_{\mathrm{tr}}$ (the height at the first and last domains) is fixed. This is explained in depth in [19].
@@
We include in Table 1 the parameters for all protocols used in this work, when applicable: domain number $n_{d}$, domain length $\ell$, sites involved in each transfer $L$, control parameter heights $v_{\mathrm{tr}}$ and $v_{\mathrm{tr}}^{\prime}$ (for the central domains in the 4-domain case), barrier height $v_{\mathrm{bar}}$, and the times for preparation ($t_{\mathrm{prep}}$), barrier preparation ($t_{\mathrm{bar}}$) and total duration ($t_{\mathrm{tr}}$).
States     $n_{d}$    $\ell$   $L$   $v_{\mathrm{tr}}$    $v_{\mathrm{tr}}^{\prime}$    $v_{\mathrm{bar}}$    $t_{\mathrm{prep}}$     $t_{\mathrm{bar}}$     $t_{\mathrm{tr}}$
@@
Table 1: Parameters for all transfer protocols used, and the states prepared with it.
@@ Appendix C Acquired phases and phase-shift protocols
$|\mathcal{P}_{k}\rangle$ $=\mathcal{N}_{\mathcal{P}}\left[\sum_{j=j_{0}-\ell/2}^{j_{0}-1}\left(-\frac{w}{v_{l}}\right)^{j-j_{0}+1}|j,\alpha(k)\rangle,\right.$
$\left.-\sum_{j=j_{0}+1}^{j_{0}+\ell/2}\left(-\frac{w}{v_{r}}\right)^{j_{0}-j+1}|j,\alpha(k)\rangle,\right]$ (17)
$|\mathcal{R}\rangle$ $=-\mathcal{N}_{\mathcal{R}}\sum_{j=N_{c}-\ell/2}^{N_{c}}\left(-\frac{w}{v}\right)^{-j+N_{c}+1}|j,\alpha(N)\rangle,$ (18)
@@
Appendix D Intermediate states for each protocol
For completeness, we include in this section the intermediate states of the system during each of the protocols.
The states at each step in the $\Phi$ state preparation protocol are:
$|\phi_{0}\rangle=|10\varnothing\rangle_{LR\gamma}$
$\bigg{\downarrow}(\Pi/2)^{(L)}_{01}$
$|\phi_{1}\rangle=\frac{1}{\sqrt{2}}[|10\varnothing\rangle-i|00\mathcal{L}\rangle]$
$\bigg{\downarrow}T_{\mathcal{L}\mathcal{R}}$
$|\phi_{2}\rangle=\frac{1}{\sqrt{2}}[|10\varnothing\rangle-i\zeta|00\mathcal{R}\rangle]$
$\bigg{\downarrow}(\Pi)^{(R)}_{01}$
$|\phi_{3}\rangle=\frac{1}{\sqrt{2}}[|10\varnothing\rangle-\zeta|01\varnothing\rangle]=|\Phi\rangle_{LR}\otimes|\varnothing\rangle_{\gamma},$
where $\zeta$ is the phase acquired in the transfer.
To prepare a $\Psi$ state, an $X_{R}$ gate must be performed on the final state of the protocol above, producing:
$|\psi\rangle=-\frac{i}{\sqrt{2}}[|11\varnothing\rangle-\zeta|00\varnothing\rangle]=|\Psi\rangle_{LR}\otimes|\varnothing\rangle_{\gamma},$
where the global phase shift is a consequence of our implementation for $X_{R}$, as mentioned in the main text.
For the W state protocol, the intermediate states are:
$|w_{0}\rangle=|100\varnothing\rangle_{LCR\gamma}$
$\bigg{\downarrow}(2\Pi/3)^{(L)}_{01}$
$|w_{1}\rangle=\frac{1}{\sqrt{3}}[|100\varnothing\rangle-i\sqrt{2}|000\mathcal{L}\rangle]$
$\bigg{\downarrow}T_{\mathcal{L}\mathcal{S}}$
$|w_{2}\rangle=\frac{1}{\sqrt{3}}[|100\varnothing\rangle-i\zeta\sqrt{2}|000\mathcal{S}\rangle]$
$\bigg{\downarrow}(\Pi/2)^{(C)}_{01}$
$|w_{3}\rangle=\frac{1}{\sqrt{3}}[|100\varnothing\rangle-\zeta|010\varnothing\rangle-i\zeta|000\mathcal{S}\rangle]$
$\bigg{\downarrow}T_{\mathcal{S}\mathcal{R}}$
$|w_{4}\rangle=\frac{1}{\sqrt{3}}[|100\varnothing\rangle-\zeta|010\varnothing\rangle-i\zeta^{2}|000\mathcal{R}\rangle]$
$\bigg{\downarrow}\Pi^{(R)}_{01}$
$|w_{5}\rangle=\frac{1}{\sqrt{3}}[|100\varnothing\rangle-\zeta|010\varnothing\rangle-\zeta^{2}|001\varnothing\rangle]=$
$=|W\rangle_{LCR}\otimes|\varnothing\rangle_{\gamma}.$
where $\zeta$ is the phase acquired in each transfer.
And finally, for the GHZ state protocol:
$|g_{0}\rangle=\frac{1}{\sqrt{2}}[|011\varnothing\rangle_{LCR\gamma}+\eta|211\varnothing\rangle_{LCR\gamma}]$
$\bigg{\downarrow}\Pi^{(L)}_{12}$
$|g_{1}\rangle=\frac{1}{\sqrt{2}}[|011\varnothing\rangle-i\eta|111\mathcal{L}\rangle]$
$\bigg{\downarrow}T_{\mathcal{L}\mathcal{S}}$
$|g_{2}\rangle=\frac{1}{\sqrt{2}}[|011\varnothing\rangle-i\eta\zeta|111\mathcal{S}\rangle]$
$\bigg{\downarrow}\Pi^{(C)}_{12}$
$|g_{3}\rangle=\frac{1}{\sqrt{2}}[|011\varnothing\rangle-\eta\zeta|121\varnothing\rangle]$
$\bigg{\downarrow}X_{C}$
$|g_{4}\rangle=\frac{-i}{\sqrt{2}}[|001\varnothing\rangle-\eta\zeta|121\varnothing\rangle]$
$\bigg{\downarrow}\Pi^{(C)}_{12}$
$|g_{5}\rangle=\frac{-i}{\sqrt{2}}[|001\varnothing\rangle+i\eta\zeta|111\mathcal{S}\rangle]$
$\bigg{\downarrow}T_{\mathcal{S}\mathcal{R}}$
$|g_{6}\rangle=\frac{-i}{\sqrt{2}}[|001\varnothing\rangle+i\eta\zeta^{2}|111\mathcal{R}\rangle]$
$\bigg{\downarrow}\Pi^{(R)}_{12}$
$|g_{7}\rangle=\frac{-i}{\sqrt{2}}[|001\varnothing\rangle+\eta\zeta^{2}|112\varnothing\rangle]$
$\bigg{\downarrow}X_{R}$
$|g_{8}\rangle=\frac{-1}{\sqrt{2}}[|000\varnothing\rangle+\eta\zeta^{2}|112\varnothing\rangle]$
$\bigg{\downarrow}\mathcal{A}_{R}$
$|g_{9}\rangle=\frac{i}{\sqrt{2}}[|000\varnothing\rangle+\eta\zeta^{2}|111\varnothing\rangle]=$
$=i|GHZ\rangle_{LCR}\otimes|\varnothing\rangle_{\gamma}.$
Appendix E Fidelity and bulk-edge decoupling
@@
where $|\psi_{0}\rangle$ is the intended final state, and $\rho_{Q}$ is the density matrix of the qudit system at the end of the realization, after tracing out the photonic degrees of freedom. For every protocol considered in this work, we obtain a fidelity higher than 0.99 for the pristine system, which is the current error-correction threshold [73]. We include the $\Phi$, W and GHZ states prepared as in the main text, and the $\Psi$ state, prepared as detailed in Appendix F.
@@
where $\rho_{ph}$ is the density matrix of the photonic degrees of freedom. The probability $p(\textrm{no bulk})$ is then found as $p(\textrm{no bulk})=p(\varnothing_{\gamma})+p(\textrm{edge})$. We can see that it is extremely close to 1, even in the presence of general disorder. This shows that the bulk-edge decoupling caused by the topology of the system is remarkably robust. This has been reported for other topological systems in the literature [13]. The three-qudit cases show a slightly smaller decoupling due to the use of trivial domains as barriers, which can allow some photon amplitude to leak into the bulk. This problem could be mitigated by decoupling the domains altogether by switching off some $w$ amplitudes, but we decided to keep them constant throughout in this work for simplicity.
@@
Appendix F GHZ-like protocol for Bell states
@@
Appendix G Generalization to any number of qubits
@@
where $\zeta_{k,k^{\prime}}$ is the phase acquired in the transfer from the cavity housing qubit $k$ to the cavity housing qubit $k^{\prime}$, which can be obtained using Eq. (19).
@@
The value of $\eta$ in the initial state can be chosen to fix the relative phase. Both protocols can be easily modified via the qudit-cavity coupling operations to change the relative weights between their final components, and the relative phases in the W state can be adjusted with phase gates or using the phase-shift protocols in Appendix C. We include all intermediate states in Appendix D.